OpenStudy (anonymous):

Is this equation a ellipse, circle, hyperbola or parabola: x^2-10x+25=y and why?

OpenStudy (turingtest):

x^2-10x+25=y can be factored into y=(x-5)^2 which is a parabola with vertex (5,0)

OpenStudy (anonymous):

I thought a parobla had to be in the form of someting like 1/80y^2=x

OpenStudy (anonymous):

i hope this would help http://higheredbcs.wiley.com/legacy/college/anton/0471472441/appendices/apph.pdf

OpenStudy (turingtest):

do you recognize that y=x^2 is a parabola?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

then look for something like (y-b)=(x-a)^2 Realize the similarity in forms, the only difference is that this parabola will have a vertex at (a,b). if we let a=0 and b=0, the above becomes y=x^2 which, as you know, is a parabola at vertex (0,0)

OpenStudy (anonymous):

i know that..i'm just suggesting a general method to determine the kind of conic section.. using the discriminant method, any equation of a conic section will be identified instantly..

OpenStudy (turingtest):

Oh, @kelu I thought I was talking to Cooper. But even so I think the page you linked for him is beyond most algebra students.